OpenStudy (anonymous):

solve for x: x ^{4}-13x ^{2}+36=0

OpenStudy (anonymous):

Put x^4 as t^2....And then Solve... :) Good LUck

OpenStudy (anonymous):

x ^{4}-13x ^{2}+36=0 is a quadratic equation in disguise. Very sneaky!

OpenStudy (anonymous):

Let x^2 = n Now you have n^{2}-13x n+36=0 now try it

OpenStudy (anonymous):

oh dont for get 36 = 6^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

take t=x^2 t^2 - 4t- 9t + 36 =0 t (t -4) - 9(t-4) =0 therefore t= 9,4

OpenStudy (anonymous):

Sorry let me be clear ...oh well mridul beat me to it